Given log Kow for Atrazine = 2.61, and Atrazine conc. in water = 4.2 ppb, estimate the Atrazine concentration (in ppb) in prawn tail tissue in an aquaculture tank (with no sediments). (assume system has come to equilibrium)
Answer #1

log Kow = log ( atrazine in organic layer / Atrazine in water)

2.61 = log ( X / 4.2 ppb)

( X/4.2ppb) = 10^ 2.61

X / ( 4.2 ppb) = 407.4

X = 4.2ppb x 407.4

= 1711 ppb

Thus Atrazine in ppb present in tail tissue ( organic) = 1711